40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 27 00:56:01 2013, Jenne, Update, ASS, Motivation for reactivating the ASS 
    Reply  Wed Feb 27 02:21:55 2013, Jenne, Update, ASS, Motivation for reactivating the ASS 
Message ID: 8176     Entry time: Wed Feb 27 00:56:01 2013     Reply to this: 8178
Author: Jenne 
Type: Update 
Category: ASS 
Subject: Motivation for reactivating the ASS 

I am putting a little bit of brain power into reviving the ASS, and I want to write down what the motivation is, and what the short and long-term plans are.

Why?

The IFO IR is not optimally aligned right now.  While we were at atmosphere, we should have taken the time to align the green beams to the arms until the greens were both resonating TEM00.  We were lazy, and excited to pump down, so we decided that locking on higher order modes was good enough to ensure the beams came out of vacuum.  Once we were pumped down, ITMY and ETMY were aligned to the green beam axis.  Then, the IR was aligned to this new Yarm cavity axis.  This would have been okay, and pretty close, if we had aligned the green beam all the way (used only the outside steering mirrors to resonate TEM00, after the cavity mirrors were aligned for flashing IR).  After the IR was aligned to the Ygreen axis, the rest of the IFO was aligned to this slightly-incorrect input pointing.  We want to measure the IR spot positions on ITMY and ETMY so that we can tweak the input pointing until we hit the center of both ITMY and ETMY.  Then we will align Ygreen's input pointing to this proper IR cavity axis.  The rest of the IFO alignment will also have to be redone.  This calls for a functioning A2L system, so the measurement part of the ASS.

The immediate motivation for measuring the spot positions is that the current Xarm IR axis is not at all very close to the Xgreen axis.  The other day while we were fixing up the Xend table (note in elog 8162), we found that the TRX beam to the TRX PD and the trans camera was clipped on the 2" harmonic separator (which is the first optic that the transmitted IR beam sees on the end tables).  It was clipping on the left side of the optic, if you are looking at the face of the optic.  This is the more east-erly side of the optic.  We moved that optic to the side so that we were not clipping.  Then, today when Manasa was trying to align the Xgreen beam, she found that it was clipping on the right side of the harmonic separator, the more west-erly side.  I remember seeing that the green beam was roughly centered on the harmonic separator when we were at atmosphere, so this clipping is certainly due to Yuta, Evan and my moving of the harmonic separator.  Since the end green steering optics are not very orthogonal in angle/translation, it is very difficult to translate the beam by a significant amount.  If we keep the current IR alignment, which surely isn't good anyway (you can see on the ETMXF camera that the beam isn't centered), we would probably have to move the Xgreen steering optics, which would be a total pain.  It seems that the better plan is to leave them where they are, and get the IR pointing in the correct direction, and move the harmonic separator back to where it was originally.

Short term (< few days):

Write the arm section of the existing MeasureSpotPositions.py script (in ....../scripts/ASS).  Write a wrapper script that, like ...../scripts/ASS/MC/mcassMCdecenter, calls sets up the lockins, runs MeasureSpotPositions.py, and calculates the calibrated spot positionsUse this information to hand tweak the input pointing, then realign the cavities to the new IR, and the greens to the new cavity axes.

All of the infrastructure for this is already in place in the c1ass model.   The only drawback to the current situation is the LSC output matrix only has one row for ASS, and so only one cavity can be measured at a time.  To make things faster, we could consider increasing the size of the LSC output matrix so that the 2 arms could be measured simultaneously.  This change is low priority for now.

Long term:

Make the full ASS system work. 

A major change from the current situation is that the current ASS cannot actuate on the input pointing (TT1 and TT2 for Yarm, BS for Xarm).  We want a low bandwidth servo to force the input beam to follow the cavity axis.  Implementing this will require some changes to the ASS model. 

Remeasure sensing matrix, test system.

ELOG V3.1.3-